OpenStudy (anonymous):

how to evaluate the integral: (e^sinx)cosx/(sinx+1)dx

myininaya (myininaya):

let u=sinx+1 du=cosx dx if u=sinx+1, then u-1=sinx see if this helps

myininaya (myininaya):

\[\int\limits_{}^{}\frac{e^{u-1}}{u} du\]

myininaya (myininaya):

\[\frac{1}{e}\int\limits_{}^{}\frac{e^u}{u}du\]

myininaya (myininaya):

maybe this is the wrong substitution let me look at it some more

OpenStudy (anonymous):

well, thanx, I was trying to do like u=1/(sinx+1) dv=(e^sinx)cosx du=-cosx/((sinx+1)^2) v=(e^sinx)

OpenStudy (anonymous):

but I still get stuck after this

myininaya (myininaya):

i dont think this is an elementary integral this is wht i see \[\int\limits\limits_{}^{}\frac{e^{\sin(x)}*\cos(x)}{sinx+1}dx\]

myininaya (myininaya):

is that right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I got this (e^sinx)/(sinx+1)+\[\int\limits (e^\sin(x))\cos(x)/(1+sinx)^2=\int\limits e^\sin(x)cosx/(1+sinx)\]

myininaya (myininaya):

this is an exponential integral

myininaya (myininaya):

what class is this?

OpenStudy (anonymous):

cal2

myininaya (myininaya):

what? what book?

OpenStudy (anonymous):

calculus 202

myininaya (myininaya):

the book is called calculus 202

OpenStudy (anonymous):

its not from the book

myininaya (myininaya):

this one is too hard for cal 2 i would think i haven't even heard of this exponential interal have you?

OpenStudy (anonymous):

no, I spent more than two hours on this

OpenStudy (anonymous):

its the first time I get stuck on one question like this

myininaya (myininaya):

i think this is too hard for cal 2 you should email your teach and tell him its not elementary integral

OpenStudy (anonymous):

I have already done that, but he told me this one is super duper easy, and thats why I still trying to do it

OpenStudy (anonymous):

Thanx for ur help anyway

myininaya (myininaya):

you wrote it down correctly?

myininaya (myininaya):

you have no doubts it was written incorrectly?

OpenStudy (anonymous):

yes, I look at it like thousand times

myininaya (myininaya):

it is not easy

myininaya (myininaya):

cosx is not in exponent part right?

OpenStudy (anonymous):

yes I think I might just leave it

myininaya (myininaya):

cow might say something in a second let him say something first and then give up

myininaya (myininaya):

well give up if he doesn't say anything that is positive about integrating this

OpenStudy (anonymous):

Ill see wht he say tom.

myininaya (myininaya):

i was able to write it in the form e^(e^s) and this is what one site said about this http://mathforum.org/library/drmath/view/52023.html

OpenStudy (dumbcow):

yeah i dont think this has a elementary anti-derivative

myininaya (myininaya):

and cow is really smart

myininaya (myininaya):

so i believe in him

myininaya (myininaya):

and myself i guess since he confirm what i thought lol